# Counting points on character varieties

### v1

@@ 6.3. The group $G=\mathrm{SO}_{5}$
Figure 9. Examples of the polynomial $|{Y}(\mathbb{F}_{q})|$ when $\mathfrak{g}=\mathfrak{so}_{5}$.
6.4. The group $G=G_{2}$
@@
$[G_{2},\phi_{1,0}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,3}^{\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,3}^{\prime\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{2,1}]$ $6q^{5}\Phi_{1}^{2}\Phi_{6}$ $2^{n}\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{6}q\Phi_{2}^{2}\Phi_{3}$ $2$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{2,2}]$ $2q^{5}\Phi_{1}^{2}\Phi_{3}$ $2^{n}\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{2}q\Phi_{2}^{2}\Phi_{6}$ $2$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,6}]$ $\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\Phi_{1}^{2}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{6}$ $1$ $12$ $1$ $1$ $1$
$[{A_{2}},3^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}\Phi_{1}^{2}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $1$ $1$ $6$ $1$ $3$ $2$
$[{A_{2}},2^{1}1^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{3}$ $4^{n}\Phi_{1}^{2}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q\Phi_{2}$ $2$ $6$ $1$ $3$ $2$
$[{A_{2}},1^{3}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}\Phi_{1}^{2}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q^{3}$ $1$ $6$ $1$ $3$ $2$
$[{A_{1}\times A_{1}^{\prime}},2^{1}\otimes 2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},2^{1}\otimes 1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},1^{2}\otimes 2^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},1^{2}\otimes 1^{2}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}\Phi_{1}^{2}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 6^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $3$ $1$ $-4$
$[{A_{1}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 6^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $3$ $1$ $-4$
$[{A_{1}^{\prime}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $-6^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $3$ $1$ $-2$
$[{A_{1}^{\prime}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $-6^{n}\Phi_{1}^{2}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $3$ $1$ $-2$
$[T,1^{1}]$ $q^{6}\Phi_{1}^{2}$ $12^{n}\Phi_{1}^{2}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $1$ $12$
Figure 10. The eighteen ${G}_{2}$-types.
@@
Figure 11. The twelve $\mathfrak{g}_{2}$-types.
$(g,n)$ $|{X}(\mathbb{F}_{q})|$
@@
Figure 12. Examples of the polynomial $|{X}(\mathbb{F}_{q})|$ when $G=G_{2}$.
$(g,n)$ $|{Y}(\mathbb{F}_{q})|$
@@
Figure 13. Examples of the polynomial $|{Y}(\mathbb{F}_{q})|$ when $\mathfrak{g}=\mathfrak{g}_{2}$.
6.5. Further examples
In this subsection, we give a few examples of counting polynomials obtained using our formula and the computer code [Bailey].
$G$ $|X(\mathbb{F}_{q})|$

### v2

@@ 7.6. Types for $\mathfrak{so}_{5}$ and expressions for the counting polynomial $f_{Y}$
7.7. Types for $G=G_{2}$
@@
$[G_{2},\phi_{1,0}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,3}^{\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,3}^{\prime\prime}]$ $3q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{3}q\Phi_{3}\Phi_{6}$ $1$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{2,1}]$ $6q^{5}\Phi_{1}^{2}\Phi_{6}$ $2^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{6}q\Phi_{2}^{2}\Phi_{3}$ $2$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{2,2}]$ $2q^{5}\Phi_{1}^{2}\Phi_{3}$ $2^{n}$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $\frac{1}{2}q\Phi_{2}^{2}\Phi_{6}$ $2$ $12$ $1$ $1$ $1$
$[G_{2},\phi_{1,6}]$ $\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $1$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}\Phi_{3}\Phi_{6}$ $q^{6}$ $1$ $12$ $1$ $1$ $1$
$[{A_{2}},3^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $1$ $1$ $6$ $1$ $3$ $2$
$[{A_{2}},2^{1}1^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{3}$ $4^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q\Phi_{2}$ $2$ $6$ $1$ $3$ $2$
$[{A_{2}},1^{3}]$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $2^{n}$ $q^{3}\Phi_{1}^{2}\Phi_{2}\Phi_{3}$ $q^{3}$ $1$ $6$ $1$ $3$ $2$
$[{A_{1}\times A_{1}^{\prime}},2^{1}\otimes 2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $1$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},2^{1}\otimes 1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},1^{2}\otimes 2^{1}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}\times A_{1}^{\prime}},1^{2}\otimes 1^{2}]$ $q^{4}\Phi_{1}^{2}\Phi_{2}^{2}$ $3^{n}$ $q^{2}\Phi_{1}^{2}\Phi_{2}^{2}$ $q^{2}$ $1$ $4$ $3$ $2$ $1$
$[{A_{1}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $3$ $1$ $-4$
$[{A_{1}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $-2\cdot 6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $3$ $1$ $-4$
$[{A_{1}^{\prime}},2^{1}]$ $q^{6}\Phi_{1}^{2}\Phi_{2}$ $-6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $1$ $1$ $2$ $3$ $1$ $-2$
$[{A_{1}^{\prime}},1^{2}]$ $q^{5}\Phi_{1}^{2}\Phi_{2}$ $-6^{n}$ $q\Phi_{1}^{2}\Phi_{2}$ $q$ $1$ $2$ $3$ $1$ $-2$
$[T,1^{1}]$ $q^{6}\Phi_{1}^{2}$ $12^{n}$ $\Phi_{1}^{2}$ $1$ $1$ $1$ $1$ $1$ $12$
7.8. Types for $\mathfrak{g}=\mathfrak{g}_{2}$
@@
7.9. Counting polynomials for $X$ and $Y$ when $G=G_{2}$
$(g,n)$ $f_{X}(q)$
@@ 7.9. Counting polynomials for $X$ and $Y$ when $G=G_{2}$
$(g,n)$ $f_{Y}(q)$
@@
7.10. Counting polynomials when $(g,n)=(0,3)$
$G$ $f_{X}(q)$